# PriorDiffusion: Leverage Language Prior in Diffusion Models for Monocular Depth Estimation

### v3

@@ A. Dataset Details
We train our model on two synthetic datasets, Hypersim [53] and Virtual Kitti [3], and conduct zero-shot evaluations on four additional real-world datasets that were not part of its training data, NYUv2 [60], KITTI [22], ScanNet [9], and ETH3D [59]. Details of each dataset are provided below.
@@ A.1. Training Datasets
Hypersim [53] is a photorealistic synthetic dataset designed for comprehensive indoor scene understanding, and is introduced since obtaining per-pixel ground truth labels from real images is often challenging or impossible for many essential scene understanding tasks. This dataset is created using a vast collection of synthetic scenes developed by professional artists, resulting in 77,400 images across 461 indoor scenes with detailed per-pixel annotations and corresponding ground truth geometry. HyperSim is built exclusively using publicly accessible 3D assets. It includes complete scene geometry, material properties, and lighting information for each scene. Also, it provides dense per-pixel semantic instance segmentations and comprehensive camera details for each image. Further, it decomposes each image into diffuse reflectance, diffuse illumination, and a non-diffuse residual component that captures view-dependent lighting effects. In terms of training split, as mentioned in the Experiments Section, we utilize the official dataset split to select approximately 54,000 samples from 365 scenes, ensuring that any incomplete samples are excluded. The RGB images and depth maps are resized to a resolution of 480 × 640 pixels, and the original distance measurements, defined relative to the focal point, are transformed into standard depth values relative to the focal plane.
Virtual Kitti [3, 18] Virtual KITTI is a photorealistic synthetic video dataset created for training and evaluating computer vision models on various video understanding tasks, including object detection, multi-object tracking, scene-level and instance-level semantic segmentation, optical flow, and depth estimation. The dataset comprises 50 high-resolution monocular videos (a total of 21,260 frames) generated from five distinct virtual urban environments, each presented under varying imaging and weather conditions. These virtual scenes were developed using the Unity game engine and an innovative real-to-virtual cloning technique. The synthetic videos come with precise, fully automatic annotations for 2D and 3D multi-object tracking, as well as per-pixel category, instance, flow, and depth labels. We use its upgraded version, Virtual KITTI 2 [3], which consists of the same five sequence clones as Virtual KITTI, with increased photorealism. It takes advantage of recent advancements in lighting and post-processing within the game engine, making the variations in the virtual sequences more closely mimic real-world changes in conditions. For training, we choose four scenes, comprising around 20,000 samples, and crop the images to match the resolution of the KITTI benchmark [22]. The maximum depth is capped at 80 meters.
@@ A.2. Evaluation Datasets
NYUv2 [60] dataset comprises 24,231 synchronized RGB images and depth maps at a resolution of $640\times 480$, representing various indoor scenes such as homes, offices, and commercial spaces, captured using a Microsoft Kinect. The standard split includes 249 training scenes and 215 test scenes. For our experiments, we use the official test set. Consistent with prior works [2, 73, 74, 75, 30], we exclude samples without valid ground truth, resulting in 654 valid images for evaluation. We perform evaluation on NYUv2 over a depth range spanning from $1\times 10^{-3}$ to 10 meters.
KITTI [22, 23] contains 61 driving scenes with research in autonomous driving and computer vision. It contains calibrated RGB images with synchronized point clouds from Velodyne lidar, inertial, GPS information, etc. Following prior works [2, 73, 74, 75, 30], we used Eigen split [13]. It consists of 652 testing images after filtering out images without valid ground truth. We follow the evaluation protocol of [14] for our experiments.
ScanNet [9] is an extensive RGB-D video dataset containing 2.5 million views in more than 1500 scans, annotated with 3D camera poses, surface reconstructions, and instance-level semantic segmentations. Data was collected using an RGB-D capture system (with a Kinect sensor) that includes automated surface reconstruction and crowdsourced semantic annotation. We use the same evaluation configuration of Marigold [30], where 800 images are randomly selected from the 312 official validation scenes for evaluation.
ETH3D [59] is a multi-view stereo and 3D reconstruction benchmark encompassing a diverse range of indoor and outdoor scenes. High-precision laser scanning was used to obtain the ground truth geometry. Images were captured using both a DSLR camera and a synchronized multi-camera rig with varying fields-of-view. For evaluation, following Marigold [30], we use all 454 samples that include ground truth depth maps.
@@ B.1. Visualization
When visualizing the ground truth depth map, we apply the same affine transformation we used in training. As described in the Experiments Section, we apply a linear normalization ensuring that the depth values primarily fall within the range $[-1,1]$. The affine transformation for normalization is defined as:
@@
It is important to note that, unlike Marigold [30], which applies linear fitting to the ground truth for error correction, we do not use this approach. While linear fitting can adjust predictions to more closely align with the ground truth, it does not accurately reflect the true distribution of the depth map predictions or provide a clear assessment of prediction quality. Instead, we enhance visualization by applying the same training normalization to the zero-shot evaluation dataset, avoiding linear fitting and its error correction effects, resulting in more authentic and insightful visualization outcomes.
B.2. Training details.
@@
We implemented our method using PyTorch, employing Stable Diffusion v2 [54] as the backbone and maintaining the original pre-training configuration with the $v$-objective [56]. The training process utilized the DDPM noise scheduler [26] with 1,000 diffusion steps, while at inference time, the DDIM scheduler [61] was employed with 50 sampling steps for faster results. Our training setup spanned 30,000 iterations, with an effective batch size of 32 achieved through gradient accumulation over 16 steps (with a per-step batch size of 2) to fit on a single Nvidia RTX 3090 GPU. We used the Adam optimizer with a learning rate set at $3\cdot 10^{-5}$ and included random horizontal flipping with a probability of 0.5 as data augmentation. For depth normalization, we employed a scale and shift-invariant method with clipping enabled and set the normalization range between -1.0 and 1.0, using a 0.02 min-max quantile to maintain robustness. This normalization strategy was applied during training and zero-shot evaluations for consistency. The training noise scheduler initialized from the pre-trained Stable Diffusion v2 [54] model maintained a noise strength of 0.9 and incorporated an annealed strategy to progressively reduce noise levels. We saved checkpoints every 50 iterations, with backup, validation, and visualization checkpoints set at intervals of 2,000 iterations. The training process typically converged after approximately 20,000 iterations, though we extended training to 30,000 iterations for thorough coverage. We used mean squared error (MSE) as the loss function, with reduction set to “mean” for averaged loss calculation. A customized iteration-wise exponential scheduler is applied which adjusts the learning rate iteratively using an exponential decay function. It decays the learning rate to 1% of its initial value over 25,000 iterations with a warmup phase of 100 steps. For text generation, generating a single caption for an image using LLaVA v1.6 on an RTX 3090 takes approximately 3.6 seconds, and we generate 10 captions for each image. For training, we generate 740,000 captions, using 740 GPU hours on an RTX 3090.
@@ B.3. Evaluation metric
Following the affine-invariant depth evaluation protocol [75, 51, 68, 30, 50], for each image and the predicted relative depth $y$, we fit a pair of scalars denoting the scale and shift parameters of the transformation: $(\hat{\alpha},\hat{\beta})=g_{\psi}(y,y^{*})\in\mathbb{R}^{2}$. The metric depth prediction is obtained by $\hat{y}=\hat{\alpha}\cdot y+\hat{\beta}$ such that:
@@
where $\hat{y}=\hat{\alpha}\cdot y+\hat{\beta}$ denotes the predicted metric-scale depth aligned from relative depth $y$, $(i,j)\in\Omega$ denotes an image coordinate, and $M:\Omega\mapsto\{0,1\}$ denotes a binary mask indicating valid coordinates in the ground truth depth $y^{*}$ with values greater than zero. Then, we follow [5, 39, 81, 74, 75] to evaluate using first-order threshold accuracy, calculated as:
@@ C.1. Additional visualizations
We provide additional visualization and analysis for indoor scenes in Figure 10 and outdoor scenes in Figure 11. We use several samples in the NYUv2 [60] and KITTI [22, 23] dataset across diverse types of scenes. We have provided examples in captions under each figure. These visualizations demonstrate that leveraging the language prior within the text-to-image diffusion model enhances the model’s ability to understand the geometric characteristics of the specified regions and objects. It shows that language plays a critical role in guiding the model’s attention to relevant regions and providing context for improved depth prediction. It highlights subtle or easily overlooked details, such as small objects or instances, and enhances the perception of complex scenes with multiple objects or intricate surfaces. Additionally, language descriptions offer an essential context for partially observed or occluded objects, enabling the model to infer details that visual cues alone might miss. By integrating language prior, the model achieves a more comprehensive and accurate understanding of scenes, especially in challenging scenarios.
@@
Method NYUv2 KITTI ETH3D $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ “An image” 95.7 6.1 89.8 10.7 95.1 6.8 Template A 95.8 6.0 90.3 10.6 95.5 6.4 Template B 95.7 6.1 90.5 10.7 95.6 6.5 Template C 95.9 6.0 90.2 10.6 95.3 6.6 Template D 95.6 5.8 90.4 10.5 95.4 6.4 Template E 95.8 5.9 90.3 10.7 95.6 6.7 Template F 95.9 6.0 90.5 10.5 95.4 6.6 Template G 95.8 6.1 90.3 10.6 95.5 6.3 Template H 95.7 5.9 90.2 10.7 95.7 6.3 Template I 95.8 6.1 90.6 10.5 95.4 6.6 Template J 95.8 6.0 90.5 10.6 95.6 6.4 Template K 95.7 5.8 90.4 10.5 95.5 6.3 Ours 95.9 5.9 90.6 10.4 95.7 6.5
@@
Table 3: Ablation for prompts to generate language description. Prompts are used to prompt LLaVA to generate language descriptions for each image. While the performances among different prompts may vary, they remain consistently comparable, as long as they are meaningful and mimic human descriptions.
C.2. Ablation for prompts to generate text
To study the effect of different prompts for text generation, here we use the visual question-answering model LLaVA v1.6 Mistral [40] to generate one text for each training image and testing image. The prompt we use should elicit responses that capture essential details, including the positioning of objects, their interactions, and notable features that influence monocular depth estimation. We generate different prompts using ChatGPT 4o [44], with the prompt:
We present the results in Table 3, showcasing various language descriptions generated by LLaVA under different prompts. While the performances exhibit some variation, they remain consistently comparable. This demonstrates that as long as meaningful language descriptions of 3D scenes, resembling natural descriptions provided by humans, are provided, PriorDiffusion can effectively leverage the language prior to enhance monocular depth estimation.
C.3. Different denoising steps.

### v4

@@ A. Dataset Details
We train our model on two synthetic datasets, Hypersim [roberts2021hypersim] and Virtual Kitti [VirtualKITTI], and conduct zero-shot evaluations on four additional real-world datasets that were not part of its training data, NYUv2 [NYU-Depth-V2], KITTI [KITTI], ScanNet [ScanNet], and ETH3D [ETH3D]. Details of each dataset are provided below.
@@ A.1. Training Datasets
Hypersim [roberts2021hypersim] is a photorealistic synthetic dataset designed for comprehensive indoor scene understanding, and is introduced since obtaining per-pixel ground truth labels from real images is often challenging or impossible for many essential scene understanding tasks. This dataset is created using a vast collection of synthetic scenes developed by professional artists, resulting in 77,400 images across 461 indoor scenes with detailed per-pixel annotations and corresponding ground truth geometry. HyperSim is built exclusively using publicly accessible 3D assets. It includes complete scene geometry, material properties, and lighting information for each scene. Also, it provides dense per-pixel semantic instance segmentations and comprehensive camera details for each image. Further, it decomposes each image into diffuse reflectance, diffuse illumination, and a non-diffuse residual component that captures view-dependent lighting effects. In terms of training split, for Marigold and E2E-FT, as mentioned in the Experiments Section, we utilize the official dataset split to select approximately 54,000 samples from 365 scenes, and the RGB images and depth maps are resized to a resolution of 480 × 640 pixels. For Lotus, approximately 39,000 samples are selected, and the RGB images and depth maps are resized to a resolution of 576 × 768 pixels. The original distance measurements, defined relative to the focal point, are transformed into standard depth values relative to the focal plane.
Virtual Kitti [VirtualKITTI, gaidon2016virtual] Virtual KITTI is a photorealistic synthetic video dataset created for training and evaluating computer vision models on various video understanding tasks, including object detection, multi-object tracking, scene-level and instance-level semantic segmentation, optical flow, and depth estimation. The dataset comprises 50 high-resolution monocular videos (a total of 21,260 frames) generated from five distinct virtual urban environments, each presented under varying imaging and weather conditions. These virtual scenes were developed using the Unity game engine and an innovative real-to-virtual cloning technique. The synthetic videos come with precise, fully automatic annotations for 2D and 3D multi-object tracking, as well as per-pixel category, instance, flow, and depth labels. We use its upgraded version, Virtual KITTI 2 [VirtualKITTI], which consists of the same five sequence clones as Virtual KITTI, with increased photorealism. It takes advantage of recent advancements in lighting and post-processing within the game engine, making the variations in the virtual sequences more closely mimic real-world changes in conditions. For training, we choose four scenes, comprising around 20,000 samples, and crop the images to match the resolution of the KITTI benchmark [KITTI]. The maximum depth is capped at 80 meters. For all models, the resolution is set to 352 × 1216.
@@ A.2. Evaluation Datasets
NYUv2 [NYU-Depth-V2] dataset comprises 24,231 synchronized RGB images and depth maps at a resolution of $640\times 480$, representing various indoor scenes such as homes, offices, and commercial spaces, captured using a Microsoft Kinect. The standard split includes 249 training scenes and 215 test scenes. For our experiments, we use the official test set. Consistent with prior works [bhat2021adabins, yuan2022neural, zeng2024wordepth, zeng2024rsa, Marigold], we exclude samples without valid ground truth, resulting in 654 valid images for evaluation. We perform evaluation on NYUv2 over a depth range spanning from $1\times 10^{-3}$ to 10 meters.
KITTI [KITTI, KITTI_1] contains 61 driving scenes with research in autonomous driving and computer vision. It contains calibrated RGB images with synchronized point clouds from Velodyne lidar, inertial, GPS information, etc. Following prior works [bhat2021adabins, yuan2022neural, zeng2024wordepth, zeng2024rsa, Marigold], we used Eigen split [Eigen-Split]. It consists of 652 testing images after filtering out images without valid ground truth. We follow the evaluation protocol of [Eigen-Split-Eval] for our experiments.
ScanNet [ScanNet] is an extensive RGB-D video dataset containing 2.5 million views in more than 1500 scans, annotated with 3D camera poses, surface reconstructions, and instance-level semantic segmentations. Data was collected using an RGB-D capture system (with a Kinect sensor) that includes automated surface reconstruction and crowdsourced semantic annotation. We use the same evaluation configuration of Marigold [Marigold], where 800 images are randomly selected from the 312 official validation scenes for evaluation.
ETH3D [ETH3D] is a multi-view stereo and 3D reconstruction benchmark encompassing a diverse range of indoor and outdoor scenes. High-precision laser scanning was used to obtain the ground truth geometry. Images were captured using both a DSLR camera and a synchronized multi-camera rig with varying fields-of-view. For evaluation, following Marigold [Marigold], we use all 454 samples that include ground truth depth maps.
@@ B.1. Visualization
When visualizing the ground truth depth map, we apply the same affine transformation we used in training. As described in the Experiments Section, we apply a linear normalization ensuring that the depth values primarily fall within the range [1,1][-1,1]. The affine transformation for normalization is defined as:
@@
It is important to note that, unlike Marigold [Marigold], which applies linear fitting to the ground truth for error correction, we do not use this approach. While linear fitting can adjust predictions to more closely align with the ground truth, it does not accurately reflect the true distribution of the depth map predictions or provide a clear assessment of prediction quality. Instead, we enhance visualization by applying the same training normalization to the zero-shot evaluation dataset, avoiding linear fitting and its error correction effects, resulting in more authentic and insightful visualization outcomes.
B.2. Training Details
@@
Marigold. For Marigold [Marigold], we implemented our method using PyTorch, employing Stable Diffusion v2 [StableDiffusion] as the backbone and maintaining the original pre-training configuration with the $v$-objective [salimans2022progressive]. The training process utilized the DDPM noise scheduler [DDPM] with 1,000 diffusion steps, while at inference time, the DDIM scheduler [DDIM] was employed with 50 sampling steps for faster results. Our training setup spanned 30,000 iterations, with an effective batch size of 32 achieved through gradient accumulation over 16 steps (with a per-step batch size of 2) to fit on a single Nvidia RTX 3090 GPU. We used the Adam optimizer with a learning rate set at $3\cdot 10^{-5}$ and included random horizontal flipping with a probability of 0.5 as data augmentation. For depth normalization, we employed a scale and shift-invariant method with clipping enabled and set the normalization range between -1.0 and 1.0, using a 0.02 min-max quantile to maintain robustness. This normalization strategy was applied during training and zero-shot evaluations for consistency. The training noise scheduler initialized from the pre-trained Stable Diffusion v2 [StableDiffusion] model maintained a noise strength of 0.9 and incorporated an annealed strategy to progressively reduce noise levels. We saved checkpoints every 50 iterations, with backup, validation, and visualization checkpoints set at intervals of 2,000 iterations. The training process typically converged after approximately 20,000 iterations, though we extended training to 30,000 iterations for thorough coverage. We used mean squared error (MSE) as the loss function, with reduction set to “mean” for averaged loss calculation. A customized iteration-wise exponential scheduler is applied which adjusts the learning rate iteratively using an exponential decay function. It decays the learning rate to 1% of its initial value over 25,000 iterations with a warmup phase of 100 steps. For text generation, generating a single caption for an image using LLaVA v1.6 on an RTX 3090 takes approximately 3.6 seconds, and we generate 10 captions for each image. For training, we generate 740,000 captions, using 740 GPU hours on an RTX 3090.
Lotus. For Lotus, we implemented our method using PyTorch, employing Stable Diffusion v2 [StableDiffusion] as the backbone and maintaining the original pre-training configuration. The training process utilized the DDPM noise scheduler [DDPM] with 1,000 diffusion steps (inherited from the pre-trained Stable Diffusion v2 configuration), while at inference time, the DDIM scheduler [DDIM] was employed with 1 sampling step. Our training setup spanned 20,000 iterations, with an effective batch size of 36 achieved through gradient accumulation over 3 steps (with a per-step batch size of 4) across 3 Nvidia RTX 3090 GPUs. We used the 8-bit Adam optimizer with a learning rate set at $3\cdot 10^{-5}$ and a constant learning rate scheduler without warmup. We included random horizontal flipping with a probability of 0.5 as data augmentation. For depth normalization, we employed a truncated disparity method, which was applied during training and zero-shot evaluations for consistency. We saved checkpoints every 500 iterations for Lotus-D and every 1,000 iterations for Lotus-G, with validation checkpoints set at the same intervals. The training process typically converged after approximately 20,000 iterations. The training utilized a mixed dataset strategy, combining Hypersim [roberts2021hypersim] at a resolution of $576\times 576$ (sampled with 90% probability) and VKITTI [VirtualKITTI] at a resolution of $375\times 375$ (sampled with 10% probability). For text generation, we used InternVL3-8B [zhu2025internvl3] to generate a single caption for each image. The generation process employed a specialized prompt focusing on depth estimation attributes (camera factors, scene properties, relative distances, object types, scales, illuminations, texture, visual features, occlusions, and boundaries), with a maximum token limit of 77 tokens per caption.
E2E-FT. For E2E-FT [garcia2024fine], we implemented our method using PyTorch, employing Stable Diffusion v2 [StableDiffusion] as the backbone and maintaining the original pre-training configuration with the $v$-objective [salimans2022progressive]. The training process utilized the DDPM noise scheduler [DDPM] with 1,000 diffusion steps (inherited from the pre-trained Stable Diffusion v2 configuration), while at inference time, the DDIM scheduler [DDIM] was employed with 1 sampling step. Our training setup spanned 20,000 iterations, with an effective batch size of 32 achieved through gradient accumulation over 16 steps (with a per-step batch size of 1) across 2 Nvidia RTX 3090 GPUs. We used the Adam optimizer with a learning rate set at $3\cdot 10^{-5}$ and a customized iteration-wise exponential learning rate scheduler that decays the learning rate to 1% of its initial value over 20,000 iterations with a warmup phase of 100 steps. We included random horizontal flipping with a probability of 0.5 as data augmentation. We employed mixed precision training with bfloat16 (bf16) for improved efficiency. For depth normalization, we employed a quantile-based method with clipping enabled, using the 0.02 and 0.98 quantiles to remove outliers and then normalizing the depth values to the range [-1.0, 1.0]. This normalization strategy was applied during training and zero-shot evaluations for consistency. The training utilized a mixed dataset strategy, combining Hypersim [roberts2021hypersim] (sampled with 90% probability) and VKITTI [VirtualKITTI] (sampled with 10% probability). For text generation, we used InternVL3-8B [zhu2025internvl3] to generate a single caption for each image. The generation process employed a specialized prompt focusing on depth estimation attributes (camera factors, scene properties, relative distances, object types, scales, illuminations, texture, visual features, occlusions, and boundaries), with a maximum token limit of 77 tokens per caption.
@@ B.3. Evaluation metric
Following the affine-invariant depth evaluation protocol [zeng2024rsa, DPT, DepthAnything, Marigold, midas], for each image and the predicted relative depth $y$, we fit a pair of scalars denoting the scale and shift parameters of the transformation: $(\hat{\alpha},\hat{\beta})=g_{\psi}(y,y^{*})\in\mathbb{R}^{2}$. The metric depth prediction is obtained by $\hat{y}=\hat{\alpha}\cdot y+\hat{\beta}$ such that:
@@
where $\hat{y}=\hat{\alpha}\cdot y+\hat{\beta}$ denotes the predicted metric-scale depth aligned from relative depth $y$, $(i,j)\in\Omega$ denotes an image coordinate, and $M:\Omega\mapsto\{0,1\}$ denotes a binary mask indicating valid coordinates in the ground truth depth $y^{*}$ with values greater than zero. Then, we follow [chang2021transformer, Va-depthnet, DepthCLIP, zeng2024wordepth, zeng2024rsa] to evaluate using first-order threshold accuracy, calculated as:
@@ C.1. Additional visualizations
We provide additional visualization and analysis for indoor scenes in Figure 10 and outdoor scenes in Figure 11. We use several samples in the NYUv2 [NYU-Depth-V2] and KITTI [KITTI, KITTI_1] dataset across diverse types of scenes. We have provided examples in captions under each figure. These visualizations demonstrate that leveraging the language prior within the text-to-image diffusion model enhances the model’s ability to understand the geometric characteristics of the specified regions and objects. It shows that language plays a critical role in guiding the model’s attention to relevant regions and providing context for improved depth prediction. It highlights subtle or easily overlooked details, such as small objects or instances, and enhances the perception of complex scenes with multiple objects or intricate surfaces. Additionally, language descriptions offer an essential context for partially observed or occluded objects, enabling the model to infer details that visual cues alone might miss. By integrating language prior, the model achieves a more comprehensive and accurate understanding of scenes, especially in challenging scenarios.
@@
C.2. Ablation for prompts to generate text
To study the effect of different prompts for text generation, here we use Marigold [Marigold] and the visual question-answering model LLaVA v1.6 Mistral [liu2023improved] to generate one text for each training image and testing image. The prompt we use should elicit responses that capture essential details, including the positioning of objects, their interactions, and notable features that influence monocular depth estimation. We generate different prompts using ChatGPT 4o [openai2024chatgpt4], with the prompt:
We present the results in Table 3, showcasing various language descriptions generated by LLaVA under different prompts. While the performances exhibit some variation, they remain consistently comparable. This demonstrates that diffusion-based depth estimators can be enhanced as long as they are provided with meaningful language descriptions of 3D scenes that resemble natural human descriptions.
Method NYUv2 KITTI ETH3D $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ Marigold 95.7 6.1 89.7 10.4 95.4 6.9 “An image” 95.7 6.1 89.8 10.7 95.1 6.8 Template A 95.8 6.0 90.3 10.6 95.5 6.4 Template B 95.7 6.1 90.5 10.7 95.6 6.5 Template C 95.9 6.0 90.2 10.6 95.3 6.6 Template D 95.6 5.8 90.4 10.5 95.4 6.4 Template E 95.8 5.9 90.3 10.7 95.6 6.7 Template F 95.9 6.0 90.5 10.5 95.4 6.6 Template G 95.8 6.1 90.3 10.6 95.5 6.3 Template H 95.7 5.9 90.2 10.7 95.7 6.3 Template I 95.8 6.1 90.6 10.5 95.4 6.6 Template J 95.8 6.0 90.5 10.6 95.6 6.4 Template K 95.7 5.8 90.4 10.5 95.5 6.3 Ours 95.9 5.9 90.6 10.4 95.7 6.5
@@ C.2. Ablation for prompts to generate text
Table 3: Ablation for prompts to generate language description. Prompts are used to prompt LLaVA to generate language descriptions for each image. While the performances among different prompts may vary, they remain consistently comparable, as long as they are meaningful and mimic human descriptions. Marigold in the first row is trained without text.
C.3. Different denoising steps